# Computing Balanced Solutions for Large International Kidney Exchange Schemes When Cycle Length Is Unbounded

### v1

@@ 2 Theoretical Results
Theorem 1.3 (restated). The core of every partitioned permutation game is non-empty, and it is possible to find a core allocation in polynomial time. Moreover, for partitioned permutation games of fixed width $c$, the problem of deciding if an allocation is in the core is polynomial-time solvable if $c=1$ and coNP-complete if $c\geq 2$.
@@
Due to the above, we just need to check whether there is a cycle $C$ such that $x(V(C))<|E(C)|=v(V(C))$. By the definition of $w_{x}$, such a cycle exists if and only if $w_{x}$ is not conservative (i.e. there is no negative weighted directed cycle), which can be decided in polynomial-time, for example with the Bellman-Ford algorithm.
Finally, we show that deciding if an allocation $x$ is in the core of a partitioned permutation game is coNP-complete, even if each $|V_{i}|$ has size $2$ (so $c=2$). Containment in coNP holds, as we can easily check if a coalition blocks an allocation. To prove hardness, we reduce from a special case of the NP-complete problem Exact $3$-Cover [28, 29].
{boxedminipage}
.96 Exact $3$-Cover
    Instance:
A family of 3-element subsets of $[3n]$, $\mathcal{S}=\{S_{1},\dots,S_{3n}\}$, where each element belongs to exactly three sets
Question:
Is there an exact $3$-cover for $\mathcal{S}$, that is, a subset $\mathcal{S^{\prime}}\subset\mathcal{S}$ such that each element appears in exactly one of the sets of $\mathcal{S^{\prime}}$?
@@
This gives a directed graph $G=(V,A)$. We partition $V$ in sets:
  • for each $i\in[3n]$, the set $A_{i}=\{a_{i},b_{i}\}$,
  • for each $k\in[6n]$, the set $X_{k}=\{x_{k},y_{k}\}$, and
  • for each $j\in[3n],l\in[3]$, the set $T_{j}^{l}=\{s_{j}^{l},t_{j}^{l}\}$.
@@
If we sum up all allocation values we get that
@@
Figure 2: An illustration for Theorem 1.3. The figure shows the part of the construction for a set $S_{1}=\{1,3,4\}$. Dashed arcs denote that there is a cycle through the given vertex whose vertices are not included in the picture. Dark grey bold ellipses denote the sets $V_{i}$.
We claim $I$ has an exact $3$-cover if and only if $x$ is not in the core. $\Rightarrow$” First suppose $\{S_{k_{1}},\dots,S_{k_{n}}\}$ is an exact $3$-cover in $I$. We claim that $\mathcal{P}=\{A_{i}\mid i\in[3n]\}\cup\{T_{k_{i}}^{l}\mid i\in[n],l\in[3]\}$ is a blocking coalition. We first show that $v(\mathcal{P})=12n$, which can be seen as follows. First, the $s_{j}^{l}$ and $a_{i}$ vertices in the coalition can be covered by $2$-cycles, as the corresponding sets form an exact $3$-cover. Moreover, the $b_{i}$ vertices can be covered by the $b_{i}$-cycle, as each of the $A_{i}$ countries is in $\mathcal{P}$, and finally, the $t_{j}^{l}$ vertices can be covered too, as for each $j\in\{k_{1},\dots,k_{n}\}$ all of $T_{j}^{1},T_{j}^{2},T_{j}^{3}$ belong to $\mathcal{P}$. Then, $x$ is not in the core, as
@@
$\Leftarrow$” Now suppose $x$ is not in the core. Then there is a coalition $\mathcal{P}$ with $v(\mathcal{P})>x(\mathcal{P})$. We write $\mathcal{A}=\bigcup_{i\in[3n]}A_{i}$ and $\mathcal{X}=\bigcup_{k\in{6n}}X_{k}$. We claim that $\mathcal{A}\subset\mathcal{P}$ or $\mathcal{X}\subset\mathcal{P}$. For a contradiction, suppose that neither $\mathcal{A}\subset\mathcal{P}$ nor $\mathcal{X}\subset\mathcal{P}$ holds. Clearly, $\mathcal{P}\cap(\mathcal{A}\cup\mathcal{X})\neq\emptyset$, because $m$ of the $T_{j}^{l}$ countries can only create a cycle packing of size $m$, but they each have an allocation of $11+\frac{1}{9n}>$ 1. So suppose that $|\mathcal{P}\cap(\mathcal{A}\cup\mathcal{X})|=m\leq 9n$. By our assumption, none of the participating $b_{i}$ or $x_{k}$ vertices can be covered. Hence, if there are $h\geq 1$ participating $T_{j}^{l}$ countries besides them (there must be at least one to have any cycles), then the size of the maximum cycle packing they can obtain is $h+2\min\{m,h\}\leq 2m+h$, as at best all $t_{j}^{l}$ vertices can be covered, but the other vertices can only be covered with cycles of length $2$ by pairing the $h$ $s_{j}^{l}$ vertices to $m$ $a_{i}$ or $x_{i}$ vertices. But, their assigned allocation in $x$ is at least
@@
First suppose that $\mathcal{A}\cup\mathcal{X}\subseteq\mathcal{P}$. Let the number of participating $T_{j}^{l}$ countries be $h$. Then, we have that
@@
Hence, we find that $12h>18n-1+\frac{h}{9n}>18n-$ 1, so $h>9n-1$, but it also cannot be $9n$, as $18n=18n-1+\frac{9n}{9n}$, a contradiction (as there are only $9n$ $T_{j}^{l}$ countries).
Suppose next that $\mathcal{X}\subseteq\mathcal{P}$. Let $0\leq m=|\mathcal{P}\cap\mathcal{A}|<3n$. Then, if the number of $T_{j}^{l}$ countries in $\mathcal{P}$ is $h>0$, then $v(\mathcal{P})\leq 6n+h+2\cdot h$. We can suppose that $h\leq 6n+m$, because if there are more $T_{j}^{l}$ countries, then at most $6n+m$ of their $s_{j}^{l}$ vertices can be covered, hence the remaining countries bring strictly more $x$ value than what they can increase the maximum cycle packing size with. However,
@@
Therefore, suppose that $\mathcal{A}\subseteq\mathcal{P}$, but $\mathcal{X}$ is not included in $\mathcal{P}$. Let $0\leq m=|\mathcal{P}\cap\mathcal{X}|<6n$. Now, if the number of $T_{j}^{l}$ countries in $\mathcal{P}$ is $h$, then $v(\mathcal{P})\leq 3n+h+2\cdot h$, similarly as before. Again, we can suppose that $h\leq 3n+m$ for similar reasons. Furthermore,
@@
To sum up, we showed that $\mathcal{P}$ must contain $\mathcal{A}$, must be disjoint from $\mathcal{X}$ and there must be exactly $3n$ $T_{j}^{l}$ countries inside $\mathcal{P}$. We claim that for each $j\in[3n]$, if $T_{j}^{l}\in\mathcal{P}$ for some $l\in[3n]$, then $T_{j}^{l}$ is inside $\mathcal{P}$ for all $l\in[3]$: if not then there must be at least one $t_{j}^{l}$ vertex that cannot be covered, hence $v(\mathcal{P})\leq 12n-1$, but $x(\mathcal{P})=9n-\frac{n+1}{3n}+3n+\frac{3n}{9n}>12n-1$, a contradiction. Therefore, for each set $S_{j}$, if $T_{j}^{l}\in\mathcal{P}$ for some $l\in[3]$, then $T_{j}^{l}\in\mathcal{P}$ for all $l\in[3]$, so there are exactly $n$ sets $S_{j}$, such that $T_{j}^{l}\in\mathcal{P}$.
@@
We now prove Theorem 1.4. We need some definitions and a lemma. Let $G=(V,A)$ be a directed graph with a partition $(V_{1},\ldots,V_{n})$ of $V$ for some $n\geq 1$. Recall that for a maximum cycle packing ${\mathcal{C}}$ of $G$, we let $s_{p}({\mathcal{C}})$ denote the number of arcs $(u,v)$ with $v\in V_{p}$ that belong to some directed cycle of ${\mathcal{C}}$. We say that ${\mathcal{C}}$ satisfies a set of intervals $\{I_{1},\ldots,I_{n}\}$ if $s_{p}({\mathcal{C}})\in I_{p}$ for every $p\in\{1,\ldots,n\}$.
@@
For instances $(G,{\mathcal{V}},{\mathcal{I}})$, where $G$ is a directed graph, ${\mathcal{V}}=(V_{1},\ldots,V_{n})$ is a partition of $V$ with fixed width $c$, and =${\mathcal{I}}=\{I_{1},\dots,I_{n}\}$ is a set of intervals, the problem of finding a maximum cycle packing of $G$ satisfying ${\mathcal{I}}$ is polynomially solvable if $c=1$, and NP-complete if $c\geq 2$ even if $I_{p}=[1,\infty]$ for every $p\in\{1,\ldots,n\}$ or $I_{p}=[1,1]$ for every $p\in\{1,\ldots,n\}$.
@@
The problem reduces to finding a maximum cycle packing such that each vertex in $U$ is covered. For this, we transform $G=(V,E)$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $v\in V\setminus U$ and its copy $v^{\prime}\in V^{\prime}\setminus U^{\prime}$, we add the edge $uu^{\prime}$ with weight 00 (we do not add these for the vertices of $U$). For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. It remains to find in polynomial time a maximum weight perfect matching in $H$, if there is any and check whether its weight is the same as the size of a maximum cycle packing in the original directed graph. If there is a perfect matching with that weight, then in the maximum cycle packing it corresponds to, each vertex in $U$ must be covered with a cycle. In the other direction, if there is a maximum cycle packing covering each vertex in $U$, then the perfect matching it corresponds to has the desired weight and we need no nonexistent $uu^{\prime}$ edge for any $u\in U$ indeed.
Now suppose $c\geq 2$. Containment in NP is trivial, as we can easily check if an arc set consists of vertex disjoint cycles or not, and for each county we can compute the number of incoming arcs.
@@
This gives a directed graph $G=(V,A)$. We partition $V$ in sets:
  • for each $i\in[3n]$, the set $A_{i}=\{a_{i}\}$,
  • for each $k\in[2n]$, the sets $X_{k}=\{x_{k}\},Y_{k}=\{y_{k}\}$, and
  • for each $j\in[3n],l\in[3]$, $T_{j}^{l}=\{s_{j}^{l},t_{j}^{l}\}$
@@
The interval for each set is $[1,\infty]$. Since the size of the maximum cycle packing of $G$ is $16n$, which is the same as the number of countries, if there is a solution that satisfies these intervals, then it also must satisfy the intervals $[1,1]$ for each set. Hence the last two statements of the Lemma are equivalent in this instance.
Figure 3: Illustration for Lemma 1, showing the construction for a set $S_{1}=\{1,2,4\}$. Grey bold lines mark the countries.
@@
Clearly, ${\mathcal{C}}$ is a cycle packing. Each $A_{i}$ has an incoming arc, as $S_{l_{1}},\dots,S_{l_{n}}$ was an exact $3$-cover. As there are exactly $2n$ sets not in the set cover, all of the corresponding sets $T_{j}^{l}$ have one incoming arc in a cycle of the form $\{(y_{i},x_{i}),(x_{i},t_{j_{i}}^{1}),(t_{j_{i}}^{1},t_{j_{i}}^{2}),((t_{j_{i}}^{3},y_{i})\}$, and so did each $X_{i}$ and each $Y_{i}$. Hence, all lower bounds are satisfied.
@@
From the above, we have that there are $n$ indices $j$ from $[3n]$, such that none of the sets $T_{j}^{1},T_{j}^{2},T_{j}^{3}$ have incoming arcs of this form. Hence, all these sets can only have incoming arcs from a set $A_{i}$. As each such $T_{j}^{i}$ must have one incoming arc, it follows that for all these $j$, the cycles $\{(a_{j_{1}},s_{j}^{1}),(s_{j}^{1},a_{j_{1}})\}$, $\{(a_{j_{2}},s_{j}^{2}),(s_{j}^{2},a_{j_{2}})\}$, $\{(a_{j_{3}},s_{j}^{3}),(s_{j}^{3},a_{j_{3}})\}$ are included in $\mathcal{C}$, so they are vertex disjoint. Hence, the corresponding sets must form an exact $3$-cover. ∎
@@
{boxedminipage}
.96 $q$-Exact Perfect Matching
    Instance:
An undirected bipartite graph $B=(U,W;E)$, where each edge is coloured with one of $\{1,\dots,q\}$, and numbers $k_{1},\ldots,k_{q}$.
Question:
Is there a perfect matching in $B$ consisting of $k_{q}$ edges of each colour $q$?
For $q=2$, this problem is also known as Exact Perfect Matching), which, as mentioned, was introduced by Papadimitriou and Yannakakis [32] and whose complexity status is open for more than 40 years. In the remainder of this section, we will gave both a reduction to this problem and a reduction from this problem. We start with doing the former in the proof of our next result (Theorem 2.1), from which Theorem 1.5 immediately follows.
@@
For a partitioned permutation game $(N,v)$ on a directed graph $G=(V,A)$ and a target allocation $x$, it is possible to generate the set of unordered deviations factors in $|A|^{O(n)}$ time by a randomized algorithm.
@@
Let $(N,v)$ be a partitioned permutation game with $n$ players, defined on a directed $G=(V,A)$ with vertex partition $V_{1},\ldots,V_{n}$. As mentioned, we reduce from $q$-Exact Perfect Matching for an appropriate value of $q$. From $(N,v)$ and a vector $d^{\prime}=(d_{1}^{\prime},\ldots,d_{n}^{\prime})$ with $d_{p}^{\prime}\geq 0$ for every $p\in N$, we define an undirected bipartite graph $B=(U,W;E)$ with coloured edges: for each vertex $v\in V$, there is a vertex $v^{in}\in U$ and a vertex $v^{out}\in W$ and an edge $v^{in}v^{out}\in E$ that has colour $n+1$; for each arc $(u,v)\in A$, there is an edge $u^{out}v^{in}$, which will be coloured $p$ if $v\in V_{p}$. Let $k_{n+1}=|V|-v(N)$ and, for $p\in\{1,\dots,n\}$, let $k_{p}=d_{p}^{\prime}$.
We observe that $G$ has a maximum cycle packing ${\mathcal{C}}$ with $s_{p}({\mathcal{C}})=d_{p}^{\prime}$ if and only if $B$ has a perfect matching with $k_{p}$ edges of each colour $p\in\{0,\ldots,n\}$.
As each $k_{i}$ can only have a value between 00 and $|E|=|A|+n=O(|A|)$, the above reduction implies that the set of unordered deviation vectors has size $|A|^{O(n)}$ for any allocation $x$ for $(N,v)$. We can find each of these vectors in $|A|^{O(n)}$ time by a randomized algorithm, as $q$-Exact Perfect Matching is solvable in $|E|^{\mathcal{O}(q)}$ time with $q$ colours with a randomized algorithm [26]. ∎
We cannot hope to derandomize the algorithm from Theorem 2.1 without first solving $2$-Exact Perfect Matching problem in polynomial time. In order to see this, let $B=(U,W;E)$ be a bipartite graph with $|U|=|W|=n$ for some integer $n$ whose edges in $E$ are coloured either red or blue. We construct a digraph $D$ by replacing every edge $e=uv$ with a directed $3$-cycle on arcs $(u,w_{e})$, $(w_{e},v)$, $(v,u)$, where $w_{e}$ is a new vertex that has only $u$ and $v$ as its neighbours in $D$. Let $V_{1}$ consist of all vertices $w_{e}$, for which $e$ is a red edge in $E$. Let $V_{2}=V(D)\setminus V_{1}$. Now, $(B,k,n-k)$ is a yes-instance of $2$-Exact Perfect Matching if and only if $D$ has a cycle packing ${\mathcal{C}}$ with $s_{1}({\mathcal{C}})=k$ and $s_{2}({\mathcal{C}})=3n-k$.

### v2

@@ 2 Theoretical Results
Theorem 1.3 (restated). The core of every partitioned permutation game is nonempty, and it is possible to find a core allocation in polynomial time. Moreover, for partitioned permutation games of fixed width $c$, the problem of deciding if an allocation is in the core is polynomial-time solvable if $c=1$ and coNP-complete if $c\geq 2$.
@@
So we just need to check whether there is a cycle $C$ such that $x(V(C))<|E(C)|=v(V(C))$. By the definition of $w_{x}$, such a cycle exists if and only if $w_{x}$ is not conservative (where conservative means no negative weighted cycles exist). The latter can be decided in polynomial-time, for example with the Bellman-Ford algorithm.
Finally, we show that deciding if an allocation $x$ is in the core of a partitioned permutation game is coNP-complete, even if each $|V_{i}|$ has size $2$ (so $c=2$). Containment in coNP holds, as we can check in polynomial time if a coalition blocks an allocation. To prove hardness, we reduce from a special case of the NP-complete problem Exact $3$-Cover [29, 30].
Exact $3$-Cover
    Instance: A family of 3-element subsets of $[3n]$, $\mathcal{S}=\{S_{1},\dots,S_{3n}\}$, where each element belongs to exactly three sets Question: Is there an exact $3$-cover for $\mathcal{S}$, that is, a subset $\mathcal{S^{\prime}}\subset\mathcal{S}$ such that each element appears in exactly one of the sets of $\mathcal{S^{\prime}}$?
@@
This gives a directed graph $G=(V,A)$. The players with their corresponding partition of the vertices are:
  • for each $i\in[3n]$, we have a player $A_{i}=\{a_{i},b_{i}\}$,
  • for each $k\in[6n]$, we have a player $X_{k}=\{x_{k},y_{k}\}$, and
  • for each $j\in[3n],l\in[3]$, we have a player $T_{j}^{l}=\{s_{j}^{l},t_{j}^{l}\}$.
@@
If we sum up all allocation values we find that
@@
Figure 2: An illustration for Theorem 1.3. The figure shows the part of the construction for a set $S_{1}=\{1,3,4\}$. Dashed arcs denote that there is a cycle through the given vertex whose vertices are not included in the picture. Dark grey bold ellipses denote the players.
We claim $I$ has an exact $3$-cover if and only if $x$ is not in the core.
$\Rightarrow$” First suppose $\{S_{k_{1}},\dots,S_{k_{n}}\}$ is an exact $3$-cover in $I$. We claim that $\mathcal{P}=\{A_{i}\mid i\in[3n]\}\cup\{T_{k_{i}}^{l}\mid i\in[n],l\in[3]\}$ is a blocking coalition. We first show that $v(\mathcal{P})=12n$, which can be seen as follows. First, the $s_{j}^{l}$ and $a_{i}$ vertices in the coalition can be covered by $2$-cycles, as the corresponding sets form an exact $3$-cover. Moreover, the $b_{i}$ vertices can be covered by the $b_{i}$-cycle, as each of the $A_{i}$ players is in $\mathcal{P}$, and finally, the $t_{j}^{l}$ vertices can be covered too, as for each $j\in\{k_{1},\dots,k_{n}\}$ all of $T_{j}^{1},T_{j}^{2},T_{j}^{3}$ belong to $\mathcal{P}$. Then, $x$ is not in the core, as
@@
$\Leftarrow$” Suppose $x$ is not in the core. Then there is a coalition $\mathcal{P}$ with $v(\mathcal{P})>x(\mathcal{P})$. We write $\mathcal{A}=\bigcup_{i\in[3n]}A_{i}$ and $\mathcal{X}=\bigcup_{k\in{6n}}X_{k}$. We claim that $\mathcal{A}\subset\mathcal{P}$ or $\mathcal{X}\subset\mathcal{P}$. For a contradiction, suppose that neither $\mathcal{A}\subset\mathcal{P}$ nor $\mathcal{X}\subset\mathcal{P}$ holds. Clearly, $\mathcal{P}\cap(\mathcal{A}\cup\mathcal{X})\neq\emptyset$, because $m$ of the $T_{j}^{l}$ players can only create a cycle packing of size $h$, but $h$ of them have together an allocation of $h(1+\frac{1}{9n})>h$. So suppose that $|\mathcal{P}\cap(\mathcal{A}\cup\mathcal{X})|=m\leq 9n$. By our assumption, none of the vertices $b_{i}$ or $x_{k}$ vertices can be covered. Hence, if there are $h\geq 1$ participating $T_{j}^{l}$ players besides them (there must be at least one to have any cycles), then the size of the maximum cycle packing they can obtain is $h+2\min\{m,h\}\leq 2m+h$, as at best all $t_{j}^{l}$ vertices can be covered, but the other vertices can only be covered with cycles of length $2$ by pairing the $h$ $s_{j}^{l}$ vertices to $m$ $a_{i}$ or $x_{i}$ vertices. But, their assigned allocation in $x$ is at least
@@
First suppose that $\mathcal{A}\cup\mathcal{X}\subseteq\mathcal{P}$. Let the number of participating $T_{j}^{l}$ players be $h$. Then, we have that
@@
Hence, we find that $12h>18n-1+\frac{h}{9n}>18n-$ 1, so $h>9n-1$, but it also cannot be $9n$, as $18n=18n-1+\frac{9n}{9n}$, a contradiction (as there are only $9n$ $T_{j}^{l}$ players).
Suppose next that $\mathcal{X}\subseteq\mathcal{P}$. Let $0\leq m=|\mathcal{P}\cap\mathcal{A}|<3n$. Then, if the number of $T_{j}^{l}$ players in $\mathcal{P}$ is $h>0$, then $v(\mathcal{P})\leq 6n+h+2\cdot h$. We can suppose that $h\leq 6n+m$, because if there are more $T_{j}^{l}$ players, then at most $6n+m$ of their $s_{j}^{l}$ vertices can be covered, hence the remaining players bring strictly more $x$ value than what they can increase the maximum cycle packing size with. However,
@@
Therefore, suppose that $\mathcal{A}\subseteq\mathcal{P}$, but $\mathcal{X}$ is not included in $\mathcal{P}$. Let $0\leq m=|\mathcal{P}\cap\mathcal{X}|<6n$. Now, if the number of $T_{j}^{l}$ players in $\mathcal{P}$ is $h$, then $v(\mathcal{P})\leq 3n+h+2\cdot h$, similarly as before. Again, we can suppose that $h\leq 3n+m$ for similar reasons. Furthermore,
@@
To sum up, we showed that $\mathcal{P}$ must contain $\mathcal{A}$, must be disjoint from $\mathcal{X}$ and there must be exactly $3n$ $T_{j}^{l}$ players inside $\mathcal{P}$.
We claim that for each $j\in[3n]$, if $T_{j}^{l}\in\mathcal{P}$ for some $l\in[3n]$, then $T_{j}^{l}$ is inside $\mathcal{P}$ for all $l\in[3]$: if not then there must be at least one $t_{j}^{l}$ vertex that cannot be covered, hence $v(\mathcal{P})\leq 12n-1$, but $x(\mathcal{P})=9n-\frac{n+1}{3n}+3n+\frac{3n}{9n}>12n-1$, a contradiction. Therefore, for each set $S_{j}$, if $T_{j}^{l}\in\mathcal{P}$ for some $l\in[3]$, then $T_{j}^{l}\in\mathcal{P}$ for all $l\in[3]$, so there are exactly $n$ sets $S_{j}$, such that $T_{j}^{l}\in\mathcal{P}$.
@@
Before we prove Theorem 1.4, we need some definitions and a lemma. Let $G=(V,A)$ be a directed graph with a partition $(V_{1},\ldots,V_{n})$ of $V$ for some $n\geq 1$. Recall that for a maximum cycle packing ${\mathcal{C}}$ of $G$, we let $s_{p}({\mathcal{C}})$ denote the number of arcs $(u,v)$ with $v\in V_{p}$ that belong to some directed cycle of ${\mathcal{C}}$. We say that ${\mathcal{C}}$ satisfies a set of intervals $\{I_{1},\ldots,I_{n}\}$ if $s_{p}({\mathcal{C}})\in I_{p}$ for every $p\in\{1,\ldots,n\}$.
@@
For instances $(G,{\mathcal{V}},{\mathcal{I}})$, where $G=(V,A)$ is a directed graph, ${\mathcal{V}}=(V_{1},\ldots,V_{n})$ is a partition of $V$ with fixed width $c$, and =${\mathcal{I}}=\{I_{1},\dots,I_{n}\}$ is a set of intervals, the problem of finding a maximum cycle packing of $G$ satisfying ${\mathcal{I}}$ is polynomially solvable if $c=1$, and NP-complete if $c\geq 2$ even if $I_{p}=[1,\infty]$ for every $p\in\{1,\ldots,n\}$ or $I_{p}=[1,1]$ for every $p\in\{1,\ldots,n\}$.
@@
The problem reduces to finding a maximum cycle packing such that each vertex in $U$ is covered. For this, we transform $G=(V,A)$ into a bipartite graph $H$ with partition classes $V$ and $V^{\prime}$, where $V^{\prime}$ is a copy of $V$. For each $v\in V\setminus U$ and its copy $v^{\prime}\in V^{\prime}\setminus U^{\prime}$, we add the edge $uu^{\prime}$ with weight 00 (we do not add these for the vertices of $U$). For each $(u,v)\in A$, we add the edge $uv$ with weight $1$. It remains to find in polynomial time a maximum weight perfect matching in $H$, if there is any and check whether its weight is the same as the size of a maximum cycle packing in the original directed graph. If there is a perfect matching with that weight, then in the maximum cycle packing it corresponds to, each vertex in $U$ must be covered with a cycle. In the other direction, if there is a maximum cycle packing covering each vertex in $U$, then the perfect matching it corresponds to has the desired weight and we need no nonexistent $uu^{\prime}$ edge for any $u\in U$ indeed.
Now suppose $c\geq 2$. Containment in NP is trivial, as we can check in polynomial time if an arc set consists of vertex disjoint cycles or not, and for each player we can compute the number of incoming arcs.
@@
This gives a directed graph $G=(V,A)$. We partition $V$ into players as follows:
  • for each $i\in[3n]$, we have a player $A_{i}=\{a_{i}\}$,
  • for each $k\in[2n]$, we have a player $X_{k}=\{x_{k}\},Y_{k}=\{y_{k}\}$, and
  • for each $j\in[3n],l\in[3]$, we have a player $T_{j}^{l}=\{s_{j}^{l},t_{j}^{l}\}$
@@
The interval for each player is $[1,\infty]$. Since the size of the maximum cycle packing of $G$ is $16n$, which is the same as the number of players, if there is a solution that satisfies these intervals, then it also must satisfy the intervals $[1,1]$ for each player. Hence the last two statements of the Lemma are equivalent in this instance.
Figure 3: Illustration for Lemma 1, showing the construction for a set $S_{1}=\{1,2,4\}$. Grey bold lines mark the players.
@@
Clearly, ${\mathcal{C}}$ is a cycle packing. Each $A_{i}$ has an incoming arc, as $S_{l_{1}},\dots,S_{l_{n}}$ was an exact $3$-cover. As there are exactly $2n$ sets not in the set cover, all of the corresponding players $T_{j}^{l}$ have one incoming arc in a cycle of the form $\{(y_{i},x_{i}),(x_{i},t_{j_{i}}^{1}),(t_{j_{i}}^{1},t_{j_{i}}^{2}),((t_{j_{i}}^{3},y_{i})\}$, and so did each $X_{i}$ and each $Y_{i}$. Hence, all lower bounds are satisfied.
@@
From the above, we have that there are $n$ indices $j$ from $[3n]$, such that none of the players $T_{j}^{1},T_{j}^{2},T_{j}^{3}$ have incoming arcs of this form. Hence, all these players can only have incoming arcs from a player $A_{i}$. As each such $T_{j}^{i}$ must have one incoming arc, it follows that for all these $j$, the cycles $\{(a_{j_{1}},s_{j}^{1}),(s_{j}^{1},a_{j_{1}})\}$, $\{(a_{j_{2}},s_{j}^{2}),(s_{j}^{2},a_{j_{2}})\}$, $\{(a_{j_{3}},s_{j}^{3}),(s_{j}^{3},a_{j_{3}})\}$ are included in $\mathcal{C}$, so they are vertex disjoint. Hence, the corresponding sets must form an exact $3$-cover. ∎
@@
$q$-Exact Perfect Matching
    Instance: An undirected bipartite graph $B=(U,W;E)$, where each edge is coloured with one of $\{1,\dots,q\}$, and numbers $k_{1},\ldots,k_{q}$. Question: Is there a perfect matching in $B$ consisting of $k_{q}$ edges of each colour $q$?
For $q=2$, this problem is also known as Exact Perfect Matching, which, as mentioned, was introduced by Papadimitriou and Yannakakis [33] and whose complexity status is open for more than 40 years. In the remainder of this section, we will give both a reduction to this problem and a reduction from this problem. We start with the former in the proof of our next result (Theorem 2.1), from which Theorem 1.5 immediately follows.
@@
For a partitioned permutation game $(N,v)$ on a directed graph $G=(V,A)$ and a target allocation $x$, it is possible to generate the set of unordered deviation vectors in $|A|^{O(n)}$ time by a randomized algorithm.
@@
Let $(N,v)$ be a partitioned permutation game with $n$ players, defined on a directed graph $G=(V,A)$ with vertex partition $V_{1},\ldots,V_{n}$. As mentioned, we reduce to $q$-Exact Perfect Matching for an appropriate value of $q$. From $(N,v)$ and a vector $d^{\prime}=(d_{1}^{\prime},\ldots,d_{n}^{\prime})$ with $d_{p}^{\prime}\geq 0$ for every $p\in N$, we define an undirected bipartite graph $B=(U,W;E)$ with coloured edges: for each vertex $v\in V$, there is a vertex $v^{in}\in U$ and a vertex $v^{out}\in W$ and an edge $v^{in}v^{out}\in E$ that has colour $n+1$; for each arc $(u,v)\in A$, there is an edge $u^{out}v^{in}$, which will be coloured $p$ if $v\in V_{p}$. Let $k_{n+1}=|V|-v(N)$ and, for $p\in\{1,\dots,n\}$, let $k_{p}=d_{p}^{\prime}$.
We observe that $G$ has a maximum cycle packing ${\mathcal{C}}$ with $s_{p}({\mathcal{C}})=d_{p}^{\prime}$ if and only if $B$ has a perfect matching with $k_{p}$ edges of each colour $p\in\{1,\ldots,n+1\}$.
As each $k_{i}$ can only have a value between 00 and $|E|=|A|+n=O(|A|)$, the above reduction implies that the set of unordered deviation vectors has size $|A|^{O(n)}$ for any allocation $x$ for $(N,v)$. We can find each of these vectors in $|A|^{O(n)}$ time by a randomized algorithm, as $q$-Exact Perfect Matching is solvable in $|E|^{O(q)}$ time with $q$ colours with a randomized algorithm [27]. ∎
We cannot hope to derandomize the algorithm from Theorem 2.1 without first solving $2$-Exact Perfect Matching problem in polynomial time, as we now show.
Theorem 2.2
Every instance of $2$-Exact Perfect Matching can be reduced in polynomial time to checking whether a partitioned permutation game $(N,v)$ with only 2 players has a solution with no deviation from a target allocation $x$.
Proof
Take an instance $I=(B,k_{1},k_{2})$ of $2$-Exact Perfect Matching. Let $B=(U,W;E)$ be the bipartite graph in $I$ with $|U|=|W|=n$ for some integer $n$ whose edges are coloured either red (colour $1$) or blue (colour $2$). We may suppose that $n=k_{1}+k_{2}$, otherwise $I$ is clearly a no-instance. We construct a digraph $G$ from $B$ by replacing every edge $e=uw$ with a directed $3$-cycle on arcs $(u,v_{e})$, $(v_{e},w)$, $(w,u)$, where $w_{e}$ is a new vertex that has only $u$ and $v$ as its neighbours in $G$.
Let the first player’s set $V_{1}$ consist of all vertices $v_{e}$, for which $e$ is a red edge in $E$ and let the other player’s vertex set be $V_{2}=V(G)\setminus V_{1}$. This defines a partitioned permutation game $(N,v)$. Let the target allocation $x$ be given by $x=(k_{1},3n-k_{1})$.
We claim that $(B,k_{1},k_{2})$ is a yes-instance of $2$-Exact Perfect Matching if and only if $G$ has a cycle packing ${\mathcal{C}}$ with $s_{1}({\mathcal{C}})=k_{1}$ and $s_{2}({\mathcal{C}})=3n-k_{1}$. Indeed, from a perfect matching $M$ of $B$ containing exactly $k_{1}$ red edges, we can subsitute each edge $uw\in M$ by the corresponding $3$-cycle $(u,v_{e},w)$, such that exactly $k_{1}$ vertices from $V_{1}$ and $3n-k_{1}$ from $V_{2}$ are covered. It is also clear that this is a maximum size cycle packing, as at most $n$ of the $v_{e}$ vertices can be covered in any cycle packing. In the other direction, suppose we have a maximum size cycle packing with $s_{1}({\mathcal{C}})=k_{1}$ and $s_{2}({\mathcal{C}})=3n-k_{1}$. Then, it covers $k_{1}$ of the vertices $v_{e}$ corresponding to red edges and $n-k_{1}=k_{2}$ of the vertices $v_{e}$ corresponding to the blue edges. Also, it must be that for each such covered vertex $v_{e}$ for $e=uw$, we have the arcs $(u,v_{e})$ and $(v_{e},w)$ in ${\mathcal{C}}$. Therefore, these edges induce a perfect matching in $B$ with exactly $k_{1}$ red and $k_{2}$ blue edges.